The capacitors have values C1 = 2.5 μF, C2 = 4.0 μF, C3 = 6.0 μF and C4 = 8.0 μF. Assume that the capacitors are connected in series. Find the equivalent capacitance of the circuit.


1
Expert's answer
2022-03-21T12:42:21-0400

Explanations & Calculations


  • If the equivalent capacitance is "\small C_e", then,

"\qquad\qquad
\begin{aligned}
\small \frac{1}{C_e}&=\small \frac{1}{2.5}+\frac{1}{4.0}+\frac{1}{6.0}+\frac{1}{8.0}\\
\small C&=\small 1.1\,\mu F
\end{aligned}"

  • When connected in series, the equivalent capacitance drops to a value even lower than the lowest capacitance in the arrangement.